# Approximation in Hölder spaces

### v2

@@ 1.2. Applications on $X=\mathbb{R}^{n}$
If $\alpha(p,q)>1,$ when $q>p,$ then $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ consists of the constant functions and the commutator is bounded iff it is the zero operator iff the symbol is constant. Therefore, regarding compactness, only the range $\alpha(p,q)\leq 1$ is interesting to study. When $\alpha(p,q)=1,$ Guo et al. [GHWY21, Theorem 1.6.] showed that the commutator is compact iff it is the zero operator, again. Now, with $p,q\in(1,\infty),$ and such that if $q>p,$ then $\alpha(p,q)<1,$ for a wide class of non-degenerate SIOs there holds that
$[b,T]\in\mathcal{K}\left(L^{p}(\mathbb{R}^{n}),L^{q}(\mathbb{R}^{n})\right)\Longleftrightarrow b\in Y^{p,q}(\mathbb{R}^{n}):=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{X^{p,q}(\mathbb{R}^{n})}.$ (6)
Emphasizing the recognition of the correct conditions on the symbol, (6) is due to Uchiyama [Uch1978], when $p=q$; due to Guo, He, Wu and Yang [GHWY21], when $q>p;$ and due to Hytönen, Li, Tao and Yang [HLTY2023], when $q<p.$ We remark that it was noted recently [Oik24] that in (6) the sufficiencies of $Y^{p,q}(\mathbb{R}^{n})$ for commutator compactness in all three cases ``$q<p",$`$q=p"$ and ``$q>p"$ all follow easily from each other by classical interpolation of compactness methods. To obtain also the necessity directions on the line (6), it is important to have descriptions of $Y^{p,q}(\mathbb{R}^{n}),$ in the cases $q\geq p,$ in terms of vanishing mean oscillation criteria. These criteria follow immediately from the existence of approximations, but to establish approximations from vanishing mean oscillation criteria takes more work. The existence of nice approximations beginning from vanishing type criteria is exactly the topic of the present article, remarkably. A particular corollary of our Theorem 1.4 is the following theorem.
@@
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2), (3) and (4). Then, there holds that
@@
To connect our pointwise conditions with vanishing mean oscillation type conditions and to establish the two descriptions of these spaces (one through approximations and one through vanishing mean oscillations) we next set some definitions. For simplicity, we let $X=\mathbb{R}^{n}$ be the Euclidean space, but $Y$ is still allowed to be an arbitrary Banach space.
@@
$\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n})}=\sup_{Q\in\mathcal{Q}}\mathcal{O}^{\omega}(f;Q)<\infty,\qquad\mathcal{O}^{\omega}(f;Q):=\frac{1}{\omega(\ell(Q))}\fint_{Q}|f-\langle f\rangle_{Q}|_{Y}.$
@@
$\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y):=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to 0}\sup_{\begin{subarray}{c}Q\in\mathcal{Q},\\ \ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y):=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}Q\in\mathcal{Q},\\ \ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y):=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}Q\in\mathcal{Q},\\ \operatorname{\mathit{d}}(Q,0)>\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y):=\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y)\cap\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)\cap\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$
One of the most interesting results in Guo et al. [GHWY21, Theorem 1.7.] states that $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}^{\alpha}(\mathbb{R}^{n})},$ for $\alpha\in(0,1);$ this is exactly the equivalence between approximations and vanishing mean oscillations mentioned above. (We also note that $\operatorname{CMO}^{\alpha}=\mathrm{VMO}^{\alpha}$ for Guo et al.) Importantly for $\alpha=0,$ the approximation $\mathrm{VMO}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}(\mathbb{R}^{n})}$ was already present in the work of Uchiyama [Uch1978]. Uchiyama’s construction does not easily translate to the case $\alpha\in(0,1),$ and Guo et al. [GHWY21] provided a non-trivial new approximation that depends on the existence of a dyadic grid on the underlying space $\mathbb{R}^{n}$ (as does Uchiyama’s original approximation) and a careful construction using this grid. One fact of our result 1.8, and in particular of its proof, is that it provides a conceptually easier approximation using only basic mollification and a (somewhat) basic truncation of the support.
Finally, to make the connection with the mean oscillations of [GHWY21, Theorem 1.7] and the pointwise conditions of Theorem 1.8 we assume the following summability condition
@@
Let $Y$ be a Banach space and the modulus $\omega$ satisfy $\omega(\infty)=\infty$, (4) and (7). Then, there holds that
$*.\displaystyle\frac{\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}}{\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}}+\frac{\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}}{\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}.$ (8)
Moreover, for each $\Gamma\in\{\operatorname{small},\operatorname{far},\operatorname{large}\},$ there holds that
@@
The proof-idea of Theorem 1.11 of linking a pointwise difference with a telescoping sum of differences of averages traces back to Meyers [Mey1964], who proved a version of (8) with the modulus $t^{\alpha},$ for $\alpha>0.$ It seems to us that the condition (7) is new, but it is natural summability condition, after all. To our best knowledge, the identification (9) is new. Combining Theorems 1.8 and 1.11, we obtain the following.
@@
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2), (3), (4) and (7). Then, there holds that
@@ 2. Bounded support approximation
In this section we prove Theorem 1.4 and then show how to easily obtain from this Theorem 1.8. At the end of this section, we also prove Theorem 1.11. Recall that we simplify notation by denoting $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ We begin with two simple lemmas.
@@
First, note that if $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and has bounded support, then, by property (3) of $\omega,$ $f$ belongs to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ Since $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ is closed under limits with respect to the seminorm $\dot{C}^{0,\omega},$ the inclusion $\overline{\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap C_{\operatorname{bs}}(X,Y)}\subset\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ holds true.
@@
Notice that $f\circ\tau-f(0)$ is zero outside $B(0,2M),$ and hence $f\circ\tau$ has bounded support. (Recall that $\dot{C}^{0,\omega}$ is defined modulo constant equivalence classes.) To actually verify the claims of (19) we begin checking some properties of the function $\tau.$ First, we verify
@@
$\begin{split}\|\tau(x)-\tau(z)\|&\leq\Big{|}\left(2-\frac{\|x\|}{M}\right)^{2}-\left(2-\frac{\|z\|}{M}\right)^{2}\Big{|}\|z\|+\left(2-\frac{\|x\|}{M}\right)^{2}\|x-z\|\\ &\leq\left(2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\right)\frac{\|x-z\|}{M}\|z\|+\left(2-\frac{\|x\|}{M}\right)^{2}\|x-z\|\\ &\leq\left(4-\frac{2M}{M}\right)\frac{\|x-z\|}{M}(2M)+\left(2-\frac{M}{M}\right)^{2}\|x-z\|=5\|x-z\|,\end{split}$ (21)
where in the second bound we used the basic identity $a^{2}-b^{2}=(a-b)(a+b).$ As $\tau$ is obviously $1$-Lipschitz in the sets $B(0,M)$ and $B(0,2M)^{c},$ and $\tau$ is continuous in $X,$ we deduce (20) for all $x,z\in X.$ Then we show the following contraction property:
@@
$\|\tau(x)-\tau(z)\|$ $\leq\left(2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\right)\frac{\|x-z\|}{M}\|z\|+\left(2-\frac{\|x\|}{M}\right)^{2}\|x-z\|$
$\leq\left(4\sqrt{\frac{R}{M}}+\frac{R}{M}\right)\|x-z\|\leq 5R\,\theta_{2}(M)\|x-z\|.$
@@
$\begin{split}\left|g(x)-g(z)\right|&=\left|f(\tau(x))-f(\tau(z))\right|\\ &\leq\left(\sup_{u\in B(0,R)^{c}}\sup_{v\in X}\frac{\left|f(u)-f(v)\right|}{\omega(\|u-v\|)}\right)\omega(\|\tau(x)-\tau(z)\|)\lesssim_{C_{\operatorname{db}}}\varepsilon\omega(\|x-z\|).\end{split}$ (29)
The case $\tau(z)\in B(0,R)^{c}$ is symmetrical. Now consider $\tau(x),\tau(z)\in B(0,R)$ and we split into subcases. But first observe that for those $x,z$ with $\|x-z\|\geq M^{1/4},$ using that $f\in\dot{C}^{0,\omega}$ we can write
@@
Provided that $M=M(\varepsilon,R)$ is large enough, the condition $\omega(\infty)=\infty$ says that the last term can be made smaller than $\varepsilon\omega(M^{1/4})\leq\varepsilon\omega(\|x-z\|),$ and thus (28) holds for these $x$ and $z.$ Therefore, during the rest of the verification of (28) we can assume that $\|x-z\|\leq M^{1/4},$ whenever needed. Notice also that in the case $x,z\in X\setminus B(0,2M)$ there is nothing to prove, as then $g(x)=g(z)=0.$ Thus, we can assume that both $x,z\in B(0,10M).$ In this case we bound
@@
Chaining the estimates (30) and (31), we have shown (28) for $x,z$ as in the present case. Finally we consider the case $\tau(x),\tau(z)\in B(0,R),$ and $x\in B(0,2M)\setminus B(0,M)$ and $z\in X\setminus B(0,10M).$ We pick an auxiliary point $y$ such that $|y|=2M.$ Using $\tau(z)=\tau(y)=0,$ the bound (31) (valid by $x,y\in B(0,10M)$) and that $\omega$ is non-decreasing we obtain
@@
which is a bound of the correct form. The case $\tau(x),\tau(z)\in B(0,R)$ and $z\in B(0,2M)$ and $x\in X\setminus B(0,10M)$ is symmetrical to the last one. We now fix the parameters $M\gg R\gg r$ so that both (27) and (28) hold. Then, we show the right claim on the line (19). If $x,z\in B(0,M)$ there is nothing to prove, since $\tau(x)=x$ and $\tau(z)=z.$ Finally, we check that

### v3

@@ 1.3. Applications on the Euclidean space
When $\alpha(p,q)>1,$ $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ consists of the constant functions and the commutator is bounded iff it is the zero operator (hence compact, in particular). Thus: for compactness only those exponents $p,q$ that result in $\alpha(p,q)\leq 1$ are interesting to study. When $\alpha(p,q)=1$ it is the case [GHWY21, Theorem 1.7. & 1.8.] that the commutator is compact iff it is the zero operator, again. Now for all $p,q\in(1,\infty)$ such that $\alpha(p,q)<1,$ there holds that
$[b,T]\in\mathcal{K}\left(L^{p}(\mathbb{R}^{n}),L^{q}(\mathbb{R}^{n})\right)\Leftrightarrow b\in Y^{p,q}(\mathbb{R}^{n}):=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{X^{p,q}(\mathbb{R}^{n})}.$ (6)
The characterization (6) is due to Uchiyama [Uch1978], when $p=q$; due to Guo, He, Wu and Yang [GHWY21], when $q>p;$ and due to Hytönen, Li, Tao and Yang [HLTY2023], when $q<p.$ When $q>p$ and $\alpha(p,q)\in(0,1),$ a result of Meyer’s [Mey1964] states that $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n})=\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n}).$ Thus, the following Theorem 1.6 provides, in particular, a new characterization of the space $Y^{p,q}(\mathbb{R}^{n})$ in (6).
@@
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2), (3) and (4). Then,
@@
To obtain the implications ``"``"`$\Rightarrow$"` ` ⇒ " in (6) it is very useful to have a description of $Y^{p,q}(\mathbb{R}^{n}),$ in the cases $q\geq p,$ in terms of some vanishing mean oscillation (VMO) criteria. Such VMO criteria (Definition 1.8, below) usually follow immediately from the existence of approximations, but to establish approximations from VMO criteria is delicate. Recall that in the previous section we established existence of approximations beginning from some pointwise vanishing criteria $\dot{\operatorname{VC}}.$ Below, as Theorem 1.9, we provide the correspondence between these $\mathrm{VMO}$ and the $\dot{\operatorname{VC}}$ type criterions, which allows us, when $X=\mathbb{R}^{n},$ to reformulate all of our results, among them Theorem 1.6, in terms of mean oscillations.
1.4. Formulation as mean oscillation conditions
@@
$\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n})}=\sup_{Q}\mathcal{O}^{\omega}(f;Q)<\infty,\qquad\mathcal{O}^{\omega}(f;Q)=\frac{1}{\omega(\ell(Q))}\fint_{Q}|f-\langle f\rangle_{Q}|_{Y},$
where the supremum is taken over all cubes in $\mathbb{R}^{n}.$
@@
$\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to 0}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\operatorname{\mathit{d}}(Q,0)>\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
where the suprema are taken over all cubes. Then, define
$\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y)$ $=\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y)\cap\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)\cap\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$
We continue by noting that the approximation $\mathrm{VMO}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}(\mathbb{R}^{n})}$ was already present in the classical work of Uchiyama [Uch1978]. Uchiyama’s construction does not easily translate to the case $\alpha\in(0,1)$ and thus in [GHWY21] a non-trivial approximation for $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}^{\alpha}(\mathbb{R}^{n})}$ was provided that relies, more or less, on the existence of dyadic grids on the underlying space $\mathbb{R}^{n}$ (as does Uchiyama’s original approximation) and a careful construction using these grids. Such a construction, when $X$ is an arbitrary normed space, has no hope of working, simply because there is no grid available. Therefore, one more fact of our Theorem 1.6, in particular of its proof, is that it provides a conceptually easier approximation that uses only basic mollification and a carefully chosen truncation of the support. To make the connection with the mean oscillations and the pointwise conditions, we assume the following summability condition
@@
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (4) and (7). Then,
$\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)},\qquad\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$ (8)
Moreover, if in addition $\omega(\infty)=\infty$, then for each scale $\Gamma\in\{\operatorname{small},\operatorname{far},\operatorname{large}\},$ we have
@@
The proof-idea of Theorem 1.9 goes back to Meyers [Mey1964], who obtained a version of the bounds (8) with the modulus $t^{\alpha},$ for $\alpha>0.$ The condition (7) appears new, but we recognize that it is a natural summability condition, after all. The identification (9) is completely new. Combining Theorems 1.6 and 1.9, we immediately obtain the following.
@@
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2), (3), (4) and (7). Then,
@@
1.5. Applications to compactness of the bi-commutator
Banach space valued operators and function spaces often appear in bi-parameter harmonic analysis. Then, it is maybe no surprise that the Banach valued Theorems (1.6), 1.10 and 1.9 have already found applications to bi-parameter harmonic analysis in the recent work [MarOik24]. We turn to the details.
We now consider $\mathbb{R}^{n}=\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$ as a bi-parameter space, take non-degenerate CZOs $T_{i}$ on $\mathbb{R}^{n_{i}}$ and $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C}).$ Then, the operator of interest, the bi-commutator, is
$[T_{2},[b,T_{1}]]=T_{2}[b,T_{1}]-[b,T_{1}]T_{2}=T_{2}bT_{1}-T_{2}T_{1}b-bT_{1}T_{2}+T_{1}bT_{2},$
where $T_{i}$ acts on the variable $x_{i}.$ The product BMO upper bound for the $L^{p}\to L^{p}$ boundedness of the bi-commutator was first obtained with the Hilbert transforms in Ferguson, Sadosky [FerSad], and in full generality of completely arbitrary CZOs later by Dalenc, Ou [DaO16]. The corresponding lower bound in terms of product $\mathrm{BMO}$ for special singular integrals (Hilbert/Riesz transforms) has a somewhat complicated history. While the product $\mathrm{BMO}$ is still believed to be the characterizing condition, the existing proofs have been reported to have a gap, see e.g. [Ferguson2002, HTV2021ce, AHLMO2021], and to the best of our knowledge a fix has not yet been published. The existence of this gap adds considerable interest to all natural questions involving the bi-commutator.
Now, in place of considering $L^{p}$-to-$L^{p}$ bounds it is natural to consider mixed $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ bounds, where we allow $q_{i}\not=p_{i},$ this was done recently in [AHLMO2021, Theorem 1.1.]. The exponents that directly connect with our work are those for which $\beta_{i}:=n_{i}(1/p_{i}-1/q_{i})\geq 0$ and $\beta_{i}>0$ for some $i=1,2,$ which we assume here and below. Then, the bi-parameter space to characterize mixed $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ boundedness turned out to be
$\|b$ $\|_{\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})}=\sup_{R=I_{1}\times I_{2}}\mathcal{O}^{\beta_{1},\beta_{2}}(b,I_{1}\times I_{2})$
$=\sup_{R=I_{1}\times I_{2}}\ell(I_{1})^{-\beta_{1}}\fint_{I_{1}}\ell(I_{2})^{-\beta_{2}}\fint_{I_{2}}|b(x_{1},x_{2})-\langle b(x_{1},\cdot)\rangle_{I_{2}}-\langle b(\cdot,x_{2})\rangle_{I_{1}}+\langle b\rangle_{I_{1}\times I_{2}}|\,\mathrm{d}x_{2}\,\mathrm{d}x_{1}.$
The space $\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})$ is in fact a $\mathrm{BMO}$-valued or $\dot{C}^{0,\alpha}$-valued Hölder space , see [AHLMO2021, Section 3]. Then, very recently [MarOik24, Theorem 1.1.] the corresponding $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ compactness was characterized through the VMO analogue $\operatorname{VMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n}).$ Moreover, as a corollary of these truly mixed norm compactness characterizations the natural product BMO condition in the non-mixed $L^{p}\to L^{p}$ case was attained as a sufficient condition. One of the crucial theorems in [MarOik24] is the following theorem, which is, in fact, a corollary of the approximations developed in this article.
Theorem 1.11.
Let $\beta_{i}\in[0,1)$ and $\beta_{i}>0$ for some $i=1,2,$ then
$\operatorname{VMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})}.$
Theorem 1.11 is a bi-parameter analogue of Theorem 1.10 (with $Y=\mathbb{C}$), where in at least one parameter we have a vanishing Hölder type behaviour. The proof of Theorem 1.11 is split between the current article and [MarOik24]. Indeed, we record here a quantitative version of Theorem 1.6 in the body-text after the proof of Theorem 1.6 as Corollary 2.5; then, assuming Corollary 2.5, Theorem 1.11 is proved in [MarOik24, Section 3.].
Acknowledgements
C. M. was supported by the grant no. 334466 of the Research Council of Norway, “Fourier Methods and Multiplicative Analysis”. T.O. was supported by the Finnish Academy of Science and Letters, and by the MICINN (Spain) grant no. PID2020-114167GB-I00.
@@ 2. Bounded support approximation
In this section we prove Theorem 1.2 and then show how to obtain from this Theorem 1.6 and then Corollary 2.5. After these, at the end of this section, we prove Theorem 1.9. Recall that we simplify notation by denoting $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ We begin with a simple remark.
Remark 2.1.
For normed spaces $X$ and $Y,$ and $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ the set $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ is closed with respect to the $\dot{C}^{0,\omega}$-seminorm, meaning that if a sequence of functions $(f_{n})_{n}\subset\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ converges to an $f\in\dot{C}^{0,\omega}(X,Y),$ then $f\in\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ as well. Indeed, for any $\varepsilon>0,$ we find $n\in\mathbb{N}$ so that $\|f-f_{n}\|_{\dot{C}^{0,\omega}}\leq\varepsilon$. Then, for any two distinct points $x,y\in X,$ we write
$\frac{|f(x)-f(y)|}{\omega(\|x-y\|)}$ $\leq\frac{|f_{n}(x)-f_{n}(y)|}{\omega(\|x-y\|)}+\frac{|(f-f_{n})(x)-(f-f_{n})(y)|}{\omega(\|x-y\|)}$
$\leq\frac{|f_{n}(x)-f_{n}(y)|}{\omega(\|x-y\|)}+\|f-f_{n}\|_{\dot{C}^{0,\omega}(X,Y)}\leq\frac{|f_{n}(x)-f_{n}(y)|}{\omega(\|x-y\|)}+\varepsilon.$
Since $f_{n}\in\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y),$ the above clearly implies $f\in\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y).$
Consequently, in the case where $Y$ is a Banach space, $(\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y),\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)})$ is a Banach space of modulo constant equivalence classes.
The following two lemmas will be very useful.
@@
First, note that if $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and has bounded support, then, by property (3) of $\omega,$ $f$ belongs to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ Since $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ is closed under limits with respect to the seminorm $\dot{C}^{0,\omega},$ as per Remark 2.1, the inclusion ``"``"`$\subset$"` ` ⊂ " holds true.
@@
Notice that $f\circ\tau-f(0)$ is zero outside $B(0,2M),$ and hence $f\circ\tau$ has bounded support. (Recall that $\dot{C}^{0,\omega}$ is defined modulo constant equivalence classes.) To actually verify the claims of (15) we begin checking some properties of the function $\tau.$
First, we verify
@@
$\begin{split}\|\tau(x)-\tau(z)\|&\leq\Big{|}\big{(}2-\frac{\|x\|}{M}\big{)}^{2}-\big{(}2-\frac{\|z\|}{M}\big{)}^{2}\Big{|}\|z\|+\big{(}2-\frac{\|x\|}{M}\big{)}^{2}\|x-z\|\\ &\leq\big{(}2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\big{)}\frac{\|x-z\|}{M}\|z\|+\big{(}2-\frac{\|x\|}{M}\big{)}^{2}\|x-z\|\\ &\leq\big{(}4-\frac{2M}{M}\big{)}\frac{\|x-z\|}{M}(2M)+\big{(}2-\frac{M}{M}\big{)}^{2}\|x-z\|=5\|x-z\|,\end{split}$ (17)
where in the second bound we used the basic identity $a^{2}-b^{2}=(a-b)(a+b).$ As $\tau$ is obviously $1$-Lipschitz in the sets $B(0,M)$ and $B(0,2M)^{c},$ and $\tau$ is continuous in $X,$ we deduce (16) for all $x,z\in X.$
Then we show the following contraction property:
@@
$\|\tau(x)-\tau(z)\|$ $\leq\Big{(}2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\Big{)}\frac{\|x-z\|}{M}\|z\|+\Big{(}2-\frac{\|x\|}{M}\Big{)}^{2}\|x-z\|$
$\leq\Big{(}4\sqrt{\frac{R}{M}}+\frac{R}{M}\Big{)}\|x-z\|\leq 5R\,\theta_{2}(M)\|x-z\|.$
@@
$\begin{split}\left|g(x)-g(z)\right|&=\left|f(\tau(x))-f(\tau(z))\right|\\ &\leq\Big{(}\sup_{u\in B(0,R)^{c}}\sup_{v\in X}\frac{\left|f(u)-f(v)\right|}{\omega(\|u-v\|)}\Big{)}\omega(\|\tau(x)-\tau(z)\|)\lesssim_{C_{\operatorname{db}}}\varepsilon\omega(\|x-z\|).\end{split}$ (23)
The case $\tau(z)\in B(0,R)^{c}$ is symmetrical.
Now consider $\tau(x),\tau(z)\in B(0,R)$ and we split into subcases. But first observe that for those $x,z$ with $\|x-z\|\geq M^{1/4},$ using that $f\in\dot{C}^{0,\omega}$ we can write
@@
Provided that $M=M(\varepsilon,R)$ is large enough, the condition $\omega(\infty)=\infty$ says that the last term can be made smaller than $\varepsilon\omega(M^{1/4})\leq\varepsilon\omega(\|x-z\|),$ and thus (22) holds for these $x$ and $z.$ Therefore, during the rest of the verification of (22) we can assume that $\|x-z\|\leq M^{1/4},$ whenever needed.
Notice also that in the case $x,z\in X\setminus B(0,2M)$ there is nothing to prove, as then $g(x)=g(z)=0.$
Thus, we can assume that both $x,z\in B(0,10M).$ In this case we bound
@@
Chaining the estimates (24) and (25), we have shown (22) for $x,z$ as in the present case.
Finally we consider the case $\tau(x),\tau(z)\in B(0,R),$ and $x\in B(0,2M)\setminus B(0,M)$ and $z\in X\setminus B(0,10M).$ We pick an auxiliary point $y$ such that $|y|=2M.$ Using $\tau(z)=\tau(y)=0,$ the bound (25) (valid by $x,y\in B(0,10M)$) and that $\omega$ is non-decreasing we obtain
@@
which is a bound of the correct form. The case $\tau(x),\tau(z)\in B(0,R)$ and $z\in B(0,2M)$ and $x\in X\setminus B(0,10M)$ is symmetrical to the last one.
We now fix the parameters $M\gg R\gg r$ so that both (21) and (22) hold. Then, we show the right claim on the line (15). If $x,z\in B(0,M)$ there is nothing to prove, since $\tau(x)=x$ and $\tau(z)=z.$ Finally, we check that